• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Subedi, F. Cusanno, R. Feuerbach

    Log entry time 22:16:01 on June12,2005

    Entry number 145966

    Followups:

    keyword=Troublshooting of right hrs y_target value

    We saw the y_target (R.gold.y) value becoming worse starting with run 4473; in run 4435 and before it was ok. So something happened between June 10 and 11. Here OK means a good y-tgt resolution for any target and in the case of waterfall target we could see two Be windows. After run 4473 through the recent run 4551 we can not see the upstream Be window of the waterfall target in the R-HRS, and the downstream window is a shoulder on the primary peak.

    This is localized to the R-HRS: the L-HRS's resolution has remained constant.

    The Q1/Q2/D/Q3 current readbacks and NMR/Hall probe measurements for the "good" and "worse" are consistent. The septum current settings and readbacks were checked, and also found to be consistent. So we do not believe it is a change in the magnetic fields.

    The beam position was found to be different between the "good" and "worse" runs. This is being investigated now.

    The effect in the R-HRS is real, not an artifact of optics elements. The observed R.tr.ph (phi at the focal plane relates directly to y-tgt) is noticibly wider in the recent run 4473 than the "good" run 4435. The distributions for these C-12 runs is shown below.

    It might be that the spectrometer exit window guards for the R-HRS are still in place, but we consider that unlikely given the number of checklists that have been performed since then, and the very limited activity in the R-HRS.



    FIGURE 1

    FIGURE 2